# On the Imaginary Part of the Effective Action in de Sitter Spacetime with Different Regularization Schemes

### v2

@@ I Introduction
A central insight of quantum field theory is that the vacuum is not empty but exhibits rich physical phenomena, including fluctuations [1], entanglement [2], and, in certain backgrounds, spontaneous particle production [3, 4]. Understanding vacuum instability and particle production is essential for quantum field theory in external fields and curved spacetimes, with important applications ranging from the Schwinger effect in strong electromagnetic fields [5, 6] to cosmology in expanding universes [7]. A general and unifying framework for describing these effects is provided by the effective action, whose imaginary part quantifies the instability of the vacuum and directly signals particle production [8, 9].
@@
Two standard approaches are commonly used to extract the imaginary part of the effective action: the Bogoliubov method and the Green’s function method. The Bogoliubov method, based on canonical quantization, expresses the vacuum persistence amplitude in terms of Bogoliubov coefficients that relate the In- and Out-modes of the field. In this framework one finds schematically [9]
@@
In the path-integral based Green’s function method, the imaginary part is obtained from the Feynman Green’s function in the coincidence limit [10] (or equivalently from the heat kernel diagonal [11]). A convenient representation is [12]
@@
In flat spacetime, such as in the Schwinger effect, or in external fields acting for a finite time, the two approaches produce expressions that, while not identical in form, can be reconciled to give the same physical result [13, 14]. In de Sitter spacetime, however, the situation is more subtle: the two methods lead to apparently different expressions for the imaginary part of the effective action. This discrepancy has led to long-standing discussions about the interpretation of particle production in de Sitter space and the role of vacuum definitions [15, 16, 17, 18].
@@ II Imaginary Part of the Effective Action in de Sitter Spacetime
in which $a_{\mathbf{k}}$ and $a_{\mathbf{k}}^{\dagger}$ are respectively the annihilation and creation operators, while $u_{\mathbf{k}}(\eta)$ are the time-dependent mode functions. Different choices of mode functions specify different vacua. In particular, the in- and out-modes can be written as [10, 12]
@@
where $H^{(1)}_{i\nu}$ and $H^{(2)}_{i\nu}$ denote Hankel functions of the first and second kind, and $k=|\mathbf{k}|$ is the magnitude of the comoving momentum. The parameter $\nu=\sqrt{m^{2}-\frac{d^{2}}{4}}$ is real for $m>\frac{d}{2}$, which is the case we focus on throughout this work. The in- and out-modes, which specify the corresponding vacua in the Poincaré patch, are related by a Bogoliubov transformation,
@@
They automatically satisfy the normalization condition and exhibit the thermal Bose–Einstein structure characteristic of particle production in de Sitter spacetime  [^1] Note that this differs from the situation in global spherical coordinates, where the Bogoliubov coefficients of the in- and out-modes exhibit an explicit dependence on the spacetime dimension; in particular, for odd dimensions the in- and out-modes are identical (i.e. $\alpha=1,\beta=0$) [19, 12]. In this work, we restrict ourselves to the case of Poincaré patch.. Substituting the explicit expression of $|\alpha_{\mathbf{k}}|^{2}$ into (3), we obtain the Bogoliubov prediction for the imaginary part of the effective action as
@@
In the approach of the Green‘s function to effective action (4), the Feynman propagator $G_{F}$ is required. In the de Sitter spacetime, the in–out Feynman propagator can be constructed from the mode functions as
@@
where $\theta(x)$ denotes the Heaviside step function and $\mathcal{N}=\sqrt{1-e^{-2\pi\nu}}$ is the normalization factor. Performing the Fourier integration, one finds that the propagator depends only on the de Sitter–invariant hyperbolic distance $z=\frac{\eta^{2}+\eta^{\prime 2}-(\mathbf{x}-\mathbf{x^{\prime}})^{2}}{2\eta\eta^{\prime}}-i\epsilon$ (where $\epsilon>0$ is the infinitesimal parameter in the Feynman $i\epsilon$ prescription fixing the analytic continuation of the propagator), as expected from the maximal symmetry of the de Sitter spacetime. The result can be expressed in closed form as [12, 20]
@@
where $Q^{\mu}_{\nu}(z)$ is the associated Legendre function of the second kind. In the coincidence limit $z=1$, the in–out Feynman propagator diverges and must be regularized. Following the previous work [12], dimensional regularization together with the minimal subtraction scheme can be employed to remove the divergence. After renormalization, the finite part of the coincident propagator exhibits a characteristic dependence on the space–time dimensions: it develops a nonvanishing imaginary contribution in even dimensions, while the imaginary part vanishes in odd dimensions [12, 21], i.e.,
@@
In the above equation, we have changed the integration variable from $\bar{m}^{2}$ to $\mu=\sqrt{\bar{m}^{2}-\frac{d^{2}}{4}}$, with $\mathrm{d}\bar{m}^{2}=2\mu\mathrm{d}\mu$. The resulting integration over $\mu$ can be evaluated exactly [21].
This feature leads to a fundamental discrepancy between the Green’s function approach and the Bogoliubov approach about particle production: the existence of an imaginary part, and hence of particle production, becomes ambiguous depending on the method employed. Moreover, even in even space–time dimensions where the Green’s function approach produces a nonvanishing imaginary part, the sign of the imaginary part alternates with dimensions. As a result, the imaginary part of the effective action changes sign with dimensionality. When it turns negative, the persistence probability blows up, signaling a puzzling loss of unitarity. These striking discrepancies have led to onging discussions in the literatures, with various attempts to reconcile the two approaches [22, 17, 13, 23, 14]. To the best of our knowledge, however, no unified resolution has yet been established in which the complete expression obtained by one method is derived from the other. We observe that the integral representations underlying the two results involve fundamentally different integration variables, while they share an important structural feature: in both cases, the integrand itself is independent of the integration variable, leading to divergent expressions. These observations provide a hint towards unification: the two approaches may be reconciled within a common integral representation that incorporates all relevant variables, differing only by the regularization prescriptions applied to the resulting divergent integrals.
@@
Here we have set $\theta(0)=\frac{1}{2}$. Due to the spatial translation invariance, the integrand is independent of the spatial coordinates, so the spatial integral simply contributes an overall volume factor $V_{d}=\int\mathrm{d}^{d}\mathbf{x}$. The remaining integrations over conformal time and comoving momentum are divergent. In order to obtain a finite result, we can introduce two explicit cutoffs: a late-time cutoff $\tau\to 0^{-}$ in the time integration and a momentum cutoff $\Lambda\to+\infty$ in the comoving momentum integration, respectively. Since the expression $G_{\mathbf{k}}(\eta,\eta)$ depends only on the magnitude of the comoving momentum, it is convenient to rewrite the momentum integration in spherical coordinates. The effective action can then be re-written as
@@
The integration over $\eta$ and $k$ can be evaluated to a closed form in terms of the generalized hypergeometric function $_{2}F_{3}$. The explicit result is
@@
$\mathcal{F}(\mu,d,\lambda)=\frac{(\lambda/2)^{2i\mu}}{\Gamma(1-i\mu)^{2}}\left(\frac{\mathcal{F}_{1}(\mu,d,\lambda)}{2\mu+id}-\frac{\mathcal{F}_{2}(\mu,d,\lambda)}{2\mu}\right),$ (25)
@@
$\mathcal{F}_{2}(\mu,d,\lambda)$ $=\,_{2}F_{3}\left(\frac{1}{2}-i\mu,-i\mu;1-i\mu,1-i\mu,1-2i\mu;-\lambda^{2}\right).$ (27)
@@
where the regular phase $\varphi(\mu,d)=\arctan\left(\frac{2\mu}{d}\right)-2\arg\Gamma(1-i\mu)$ does not depend on $\lambda$. In the strict $\lambda\to 0^{+}$ limit, the second term in the integrand oscillates rapidly in $\mu$ with frequency $2\log(\lambda/2)\to+\infty$, while its amplitude is exponentially suppressed by $e^{-\pi\mu}$. By the Riemann–Lebesgue lemma [24], this term gives a vanishing contribution upon the integration over $\mu$. The expression therefore simplifies to
@@
The coefficients $a_{n}(\mu,d)$ and $b_{n}(\mu,d)$ originate from the large-argument expansion of the generalized hypergeometric functions $_{2}F_{3}$ in (25), which can in principle be constructed systematically according to Ref.[25]. These coefficients produce a hierarchy of oscillatory contributions proportional to $\lambda^{-n}\cos(2\lambda)$ and $\lambda^{-n}\sin(2\lambda)$, which represent interference effects between the late-time cutoff $\tau$ and the momentum cutoff $\Lambda$. Although they can dominate the integrand’s behavior for large but finite $\lambda$, they vanish in the distributional sense and thus do not contribute to the physical limit $\lambda\to+\infty$. From a physical perspective, we therefore retain only the smoothly varying component, represented by the non–oscillatory term proportional to $\lambda^{-d}$. Substituting this part into the integration (24), we get
@@
$\frac{1}{2\pi^{d/2}\Gamma(-\frac{d}{2})\Gamma(d+1)}=\frac{(-1)^{\tfrac{d+1}{2}}}{(4\pi)^{\frac{d+1}{2}}\Gamma\left(\frac{d+1}{2}\right)}\ ,\ d+1=2,4,6\cdots,$ (34)
@@
which correctly reproduces the imaginary part of the effective action derived from the Green’s function method (18), up to rapidly oscillating terms that will vanish in the distributional sense.
@@ III Conclusions and Discussions
  • In the opposite limit $\lambda\to+\infty$, we recover the Green’s function result, including its characteristic dependence on spacetime dimensionality (33).
@@
Our analysis should be compared with that of Akhmedov et al. [14], who attributed the discrepancy to the presence of vacuum wave functionals in the functional integral (1). In our cutoff framework, such contributions are not taken into account explicitly, however, we get consistent results in opposite limits. It seems that including the vacuum wave functionals are not very indispensable, however, on the other hand it will be interesting in future to examine whether including them can lead to a unique, ambiguity-free result. Our approach also differs from the trick used in earlier works [22, 17, 13], where a divergent momentum integral was effectively reinterpreted as a divergent time integral through scaling relations connected by physical cutoffs (see [14] for examples). While that procedure provides a way to handle divergences, it does not precisely reproduce both outcomes. By contrast, our framework unifies the two limiting results within a single regulated expression, rather than merely substituting one source of divergence for another.
@@
From a physical viewpoint, the Bogoliubov prescription has the advantage of yielding a manifestly positive imaginary part, naturally interpreted as particle production probabilities, whereas the Green’s function prescription shows a more intricate dependence on dimensionality and may even produce unphysical results such as negative probabilities. This suggests that the Bogoliubov limit corresponds more closely to a physically meaningful definition of the vacuum persistence amplitude. In the meantime, the very existence of these ambiguities signals potential limitations of the in–out formalism in genuinely non–stationary backgrounds. As emphasized by Akhmedov et al. [12, 14], the more appropriate framework in such situations is the Keldysh–Schwinger technique, which provides access to well–defined observables such as the expectation value of the energy–momentum tensor rather than ill–defined particle numbers [26].

### v3

@@ I Introduction
A central insight of quantum field theory is that the vacuum is not empty but exhibits rich physical phenomena, including fluctuations [Casimir:1948dh], entanglement [Reznik:2002fz], and, in certain backgrounds, spontaneous particle production [Hawking:1975vcx, Schwinger:1951nm]. Understanding vacuum instability and particle production is essential for quantum field theory in external fields and curved spacetimes, with important applications ranging from the Schwinger effect in strong electromagnetic fields [DUNNE_2005, Gelis_2016] to cosmology in expanding universes [Ford_2021]. A general and unifying framework for describing these effects is provided by the effective action, whose imaginary part quantifies the instability of the vacuum and directly signals particle production [DeWitt:2003pm, Parker_Toms_2009].
@@
Two standard approaches are commonly used to extract the imaginary part of the effective action: the Bogoliubov method and the Green’s function method. The Bogoliubov method, based on canonical quantization, expresses the vacuum persistence amplitude in terms of Bogoliubov coefficients that relate the In- and Out-modes of the field. In this framework one finds schematically [Parker_Toms_2009]
@@
In the path-integral based Green’s function method, the imaginary part is obtained from the Feynman Green’s function in the coincidence limit [Birrell:1982ix] (or equivalently from the heat kernel diagonal [Vassilevich_2003]). A convenient representation is [Akhmedov_2019]
@@
In flat spacetime, such as in the Schwinger effect, or in external fields acting for a finite time, the two approaches produce expressions that, while not identical in form, can be reconciled to give the same physical result [Anderson_2018, Akhmedov:2024qvi]. In de Sitter spacetime, however, the situation is more subtle: the two methods lead to apparently different expressions for the imaginary part of the effective action. This discrepancy has led to long-standing discussions about the interpretation of particle production in de Sitter space and the role of vacuum definitions [Gibbons:1977mu, kim2010, Anderson_2014, AKHMEDOV_2010].
@@ II Imaginary Part of the Effective Action in de Sitter Spacetime
in which $a_{\mathbf{k}}$ and $a_{\mathbf{k}}^{\dagger}$ are respectively the annihilation and creation operators, while $u_{\mathbf{k}}(\eta)$ are the time-dependent mode functions. Different choices of mode functions specify different vacua. In particular, the in- and out-modes can be written as [Birrell:1982ix, Akhmedov_2019]
@@
where $H^{(1)}_{i\nu}(x)$ and $H^{(2)}_{i\nu}(x)$ denote Hankel functions of the first and second kind, and $k=|\mathbf{k}|$ is the magnitude of the comoving momentum. The parameter $\nu=\sqrt{m^{2}-\frac{d^{2}}{4}}$ is real for $m>\frac{d}{2}$, which is the case we focus on throughout this work. For related discussions on the light-field case $m<\frac{d}{2}$, see Refs. [Akhmedov:2024npw, Miller:2025jbz, Jain:2025maa]. The in- and out-modes, which specify the corresponding vacua in the Poincaré patch, are related by a Bogoliubov transformation,
@@
They automatically satisfy the normalization condition and exhibit the thermal Bose–Einstein structure characteristic of particle production in de Sitter spacetime  [^1] Note that this differs from the situation in global spherical coordinates, where the Bogoliubov coefficients of the in- and out-modes exhibit an explicit dependence on the spacetime dimension; in particular, for odd dimensions the in- and out-modes are identical (i.e. $\alpha=1,\beta=0$) [Bousso_2002, Akhmedov_2019]. In this work, we restrict ourselves to the case of Poincaré patch.. Substituting the explicit expression of $|\alpha_{\mathbf{k}}|^{2}$ into (3), we obtain the Bogoliubov prediction for the imaginary part of the effective action as
@@
In the approach of the Green’s function to effective action (4), the Feynman propagator $G_{F}$ is required. In the de Sitter spacetime, the in–out Feynman propagator can be constructed from the mode functions as
@@
where $\theta(x)$ denotes the Heaviside step function and $\mathcal{N}=\sqrt{1-e^{-2\pi\nu}}$ is the normalization factor. Performing the Fourier integration, one finds that the propagator depends only on the de Sitter–invariant hyperbolic distance $z=\frac{\eta^{2}+\eta^{\prime 2}-(\mathbf{x}-\mathbf{x^{\prime}})^{2}}{2\eta\eta^{\prime}}-i\epsilon$ (where $\epsilon>0$ is the infinitesimal parameter in the Feynman $i\epsilon$ prescription fixing the analytic continuation of the propagator), as expected from the maximal symmetry of the de Sitter spacetime. The result can be expressed in closed form as [Akhmedov_2019, Fukuma_2013]
@@
where $Q^{\mu}_{\nu}(x)$ is the associated Legendre function of the second kind. In the coincidence limit $z=1$, the in–out Feynman propagator diverges and must be regularized. Following the previous work [Akhmedov_2019], dimensional regularization together with the modified minimal subtraction ($\overline{\text{MS}}$) scheme can be employed to remove the divergence. After renormalization, the finite part of the coincident propagator exhibits a characteristic dependence on the space–time dimensions: it develops a nonvanishing imaginary contribution in even dimensions, while the imaginary part vanishes in odd dimensions [Akhmedov_2019, zhou2024], i.e.,
@@
In the above equation, we have changed the integration variable from $\bar{m}^{2}$ to $\mu=\sqrt{\bar{m}^{2}-\frac{d^{2}}{4}}$, with $\mathrm{d}\bar{m}^{2}=2\mu\mathrm{d}\mu$. The resulting integration over $\mu$ can be evaluated exactly [zhou2024].
This feature leads to a fundamental discrepancy between the Green’s function approach and the Bogoliubov approach about particle production: the existence of an imaginary part, and hence of particle production, becomes ambiguous depending on the method employed. Moreover, even in even space–time dimensions where the Green’s function approach produces a nonvanishing imaginary part, the sign of the imaginary part alternates with dimensions. As a result, the imaginary part of the effective action changes sign with dimensions. When it turns negative, the persistence probability blows up, signaling a puzzling loss of unitarity. These striking discrepancies have led to onging discussions in the literatures, with various attempts to reconcile the two approaches [Mottola:1984ar, Anderson_2014, Anderson_2018, Kim:2011rx, Akhmedov:2024qvi]. To the best of our knowledge, however, no unified resolution has yet been established in which the complete expression obtained by one method is derived from the other. We observe that the integral representations underlying the two results involve fundamentally different integration variables, while they share an important structural feature: in both cases, the integrand itself is independent of the integration variable, leading to divergent expressions. These observations provide a hint towards unification: the two approaches may be reconciled within a common integral representation that incorporates all relevant variables, differing only by the regularization prescriptions applied to the resulting divergent integrals.
@@
Here we set $\theta(0)=\frac{1}{2}$ and $J_{i\nu}(x)$ denotes the Bessel function of the first kind. Due to the spatial translation invariance, the integrand is independent of the spatial coordinates, so the spatial integral simply contributes an overall volume factor $V_{d}=\int\mathrm{d}^{d}\mathbf{x}$. The remaining integrations over conformal time and comoving momentum are divergent. In order to obtain a finite result, we can introduce two explicit cutoffs: a late-time cutoff $\tau\to 0^{-}$ in the time integration and a momentum cutoff $\Lambda\to+\infty$ in the comoving momentum integration, respectively. Since the expression $G_{\mathbf{k}}(\eta,\eta)$ depends only on the magnitude of the comoving momentum, it is convenient to rewrite the momentum integration in spherical coordinates. The effective action can then be re-written as
@@
The integration over $\eta$ and $k$ can be evaluated to a closed form in terms of the generalized hypergeometric function $_{2}F_{3}$. The explicit result is given below, while the detailed steps of the derivation are presented in Appendix A,
@@
$\mathcal{F}(\mu,d,\lambda)=\frac{(\lambda/2)^{-2i\mu}}{\Gamma(1-i\mu)^{2}}\left(\frac{\mathcal{F}_{1}(\mu,d,\lambda)}{2\mu+id}-\frac{\mathcal{F}_{2}(\mu,\lambda)}{2\mu}\right),$ (25)
@@
$\mathcal{F}_{2}(\mu,\lambda)$ $=\,_{2}F_{3}\left(\frac{1}{2}-i\mu,-i\mu;1-i\mu,1-i\mu,1-2i\mu;-\lambda^{2}\right).$ (27)
@@
where the regular phase $\varphi(\mu,d)=\arctan\left(\frac{2\mu}{d}\right)-2\arg\Gamma(1-i\mu)$ does not depend on $\lambda$. In the strict $\lambda\to 0^{+}$ limit, the second term in the integrand oscillates rapidly in $\mu$ with frequency $2\log(\lambda/2)\to+\infty$, while its amplitude is exponentially suppressed by $e^{-\pi\mu}$. By the Riemann–Lebesgue lemma [Haber2018RiemannLebesgue], this term gives a vanishing contribution upon the integration over $\mu$. The expression therefore simplifies to
@@
The coefficients $a_{n}(\mu,d)$ and $b_{n}(\mu,d)$ originate from the large-argument expansion of the generalized hypergeometric functions $_{2}F_{3}$ in (25), which can in principle be constructed systematically according to Ref.[DLMF]. These coefficients produce oscillatory contributions of the form $\lambda^{-n}\cos(2\lambda)$ and $\lambda^{-n}\sin(2\lambda)$. Their appearance reflects the sensitivity of the regulated integral to the hard cutoffs. As reported in Ref.[AAD_2025], such sharp switchings generally introduce interference terms tied to the abrupt nature of the cutoff, rather than to the intrinsic dynamics of particle production. In the present calculation, the oscillatory structures exactly represent the cutoff-induced interference between the late-time cutoff at $\tau$ and the momentum cutoff at $\Lambda$. Although they can dominate the integrand’s behavior for large but finite $\lambda$, they do not yield a smooth limit as $\lambda\to+\infty$. From a physical perspective, we retain only the smoothly varying component, represented by the non–oscillatory term proportional to $\lambda^{-d}$. Substituting this part into the integration (24), we get
@@
$\frac{1}{2\pi^{d/2}\Gamma(-\frac{d}{2})\Gamma(d+1)}=\frac{(-1)^{\tfrac{d+1}{2}}}{(4\pi)^{\frac{d+1}{2}}\Gamma\left(\frac{d+1}{2}\right)}\ ,\ (d+1=2,4,6\cdots),$ (34)
@@
which correctly reproduces the imaginary part of the effective action derived from the Green’s function method (18), up to rapidly oscillating cutoff–induced terms that should not exist in the physical $\lambda\to+\infty$ limit.
@@ III Conclusions and Discussions
  • In the opposite limit $\lambda\to+\infty$, we recover the Green’s function result, including its characteristic dependence on spacetime dimension (33).
@@
Our analysis should be compared with that of Akhmedov et al. [Akhmedov:2024qvi], who attributed the discrepancy to the additional contributions arising from the vacuum wave functionals in the functional integral (1), beyond merely specifying the in–out boundary conditions. In our cutoff framework, such contributions are not taken into account explicitly, however, we get consistent results in opposite limits. It seems that including the vacuum wave functionals are not very indispensable, however, on the other hand it will be interesting in future to examine whether including them can lead to a unique, ambiguity-free result. Our approach also differs from the trick used in earlier works [Mottola:1984ar, Anderson_2014, Anderson_2018], where a divergent momentum integral was effectively reinterpreted as a divergent time integral through scaling relations connected by physical cutoffs (see [Akhmedov:2024qvi] for examples). While that procedure provides a way to handle divergences, it does not precisely reproduce both outcomes. By contrast, our framework unifies the two limiting results within a single regulated expression, rather than merely substituting one source of divergence for another.
The trade-off parameter $\lambda$ can be interpreted as the maximum physical momentum that remains accessible at late times in de Sitter spacetime (with Hubble scale $H=1$). From the structure of the integral (40 in Appendix A), the modes whose physical momenta lie below this late-time accessible scale $\lambda$ contribute to the effective action in the form as in the Green’s function approach, while the modes with physical momenta above $\lambda$ enter through the Bogoliubov-type contribution. These two classes of modes contribute to the imaginary part of the effective action through distinct integral representations, which explains why the two calculation schemes generally yield different results. If one assumes that cosmological redshift drives all physical momenta to zero at late times, the effective action reduces to the Bogoliubov result. Conversely, if one assumes that a late-time observer still has access to the full range of physical momenta, the Green’s function result will be obtained. Which prescription is physically appropriate requires careful consideration of the specific observational and theoretical circumstances. Clarifying this distinction is an interesting direction for future work.
@@
From a physical viewpoint, the Bogoliubov prescription has the advantage of yielding a manifestly positive imaginary part, naturally interpreted as particle production probabilities, whereas the Green’s function prescription shows a more intricate dependence on dimension and may even produce unphysical results such as negative probabilities. This suggests that the Bogoliubov limit corresponds more closely to a physically meaningful definition of the vacuum persistence amplitude. In the meantime, the very existence of these ambiguities signals potential limitations of the in–out formalism in genuinely non–stationary backgrounds. As emphasized by Akhmedov et al. [Akhmedov_2019, Akhmedov:2024qvi], the more appropriate framework in such situations is the Keldysh–Schwinger technique, which provides access to well–defined observables such as the expectation value of the energy–momentum tensor rather than ill–defined particle numbers [AKHMEDOV_2014LN]. Despite the fact that the Keldysh–Schwinger technique concerns a different class of physical observables, the phenomenon we report here is in fact generic in quantum field theory in de Sitter space: It is a manifestation of the well-known non-commutativity of the UV and IR limits [Polyakov:2012uc], which arises independently of the particular computational framework. When handling such divergent integrals, it is therefore important to rely on a physically motivated interpretation to select the appropriate order of limits or regularization schemes. Our unified analysis thus provides a complementary perspective, clarifying a conceptual issue that lies outside the natural scope of the Keldysh–Schwinger framework.